All Systems Operational
MKE1 Core Datacenter Infrastructure Operational
90 days ago
100.0 % uptime
Today
MKE1 Hosted Application & Dedicated Services Operational
90 days ago
99.99 % uptime
Today
Voice Over IP Services (MKE1) Operational
90 days ago
99.98 % uptime
Today
Mac Hosting (MKE1) ? Operational
90 days ago
100.0 % uptime
Today
Web & Email Hosting (MKE1) Operational
90 days ago
100.0 % uptime
Today
File Sharing (MKE1) Operational
90 days ago
100.0 % uptime
Today
PHX1 Core Datacenter Infrastructure Operational
90 days ago
100.0 % uptime
Today
PHX1 Hosted Application & Dedicated Services Operational
90 days ago
100.0 % uptime
Today
Voice Over IP Services (PHX1) Operational
90 days ago
100.0 % uptime
Today
Mac Hosting (PHX1) ? Operational
90 days ago
100.0 % uptime
Today
Customer Portal Operational
90 days ago
100.0 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
had a major outage
had a partial outage
Past Incidents
May 31, 2020

No incidents reported today.

May 30, 2020

No incidents reported.

May 29, 2020

No incidents reported.

May 28, 2020
Resolved - The cause of the packet loss has been mitigated and this incident is resolved.
May 28, 17:45 CDT
Monitoring - A fix has been implemented and our NOC is monitoring to ensure stability.
May 28, 17:31 CDT
Update - Our NOC is continuing to investigate the intermittent packet loss and is working to implement a fix.
May 28, 17:00 CDT
Identified - A small number of FreePBX Hosting dedicated servers in our MKE1 data center are experiencing some intermittent packet loss and our NOC is currently investigating.
May 28, 16:29 CDT
May 27, 2020

No incidents reported.

May 26, 2020

No incidents reported.

May 25, 2020

No incidents reported.

May 24, 2020

No incidents reported.

May 23, 2020

No incidents reported.

May 22, 2020
Resolved - This incident has been resolved. The affected fiber provider will be added back into our BGP mix at a later time after their outage has been resolved.
May 22, 16:55 CDT
Monitoring - We have removed the fiber provider from our BGP mix until the outage is resolved to ensure network stability. All traffic is being routed over our other fiber providers at this time.
May 22, 16:27 CDT
Identified - Monitoring in our MKE1 data center has detected an upstream fiber provider is currently experiencing an outage. Traffic is automatically being rerouted to our other fiber providers while we work with the provider to restore service.
May 22, 16:00 CDT
May 21, 2020

No incidents reported.

May 20, 2020

No incidents reported.

May 19, 2020

No incidents reported.

May 18, 2020

No incidents reported.

May 17, 2020

No incidents reported.